OpenStudy (anonymous):

A company makes jewelry. It costs $6,400 to make 10 pieces of jewelry and $11,200 to make 20 pieces of jewelry. Which equation models the cost, C(x), as a linear function of the number of pieces of jewelry, x? C(x) = 480x + 1600 C(x) = 1600x + 480 C(x) = 480x − 1600 C(x) = 1600x − 480

OpenStudy (ash2326):

Let cost be C(x) , x is the no. of pieces C(x)=Ax+B A and B are constants that we have to find it's given it costs $6400 to make 10 pieces and $11200 to make 20 pieces let's use the first condition 1st.....6400=10A+B and then from second condition 2nd.....11200=20A+B subtract 2nd from 1st 4800=10A A=480 substutite value of A in 1st to get B 6400=4800+B B=1600 we get C(x)=480x+1600 option A